OpenStudy (anonymous):

Seth has two similar ice-cream cones, Cone A and Cone B. The height of Cone A is 16 inches, and the height of Cone B is 4 inches. What is the ratio of the volume of Cone A to the volume of Cone B? (1 point) a. 64:1 b. 8:1 c. 4:1 d. 16:1

OpenStudy (anonymous):

16 = 4 * 4 4 = 4 * 1 So it would have to be C

OpenStudy (anonymous):

thathank you!